P(at least one head)=1−P(no head)=1−(1/2)10=1−1/1024=1023/1024.
If a fair coin is tossed 10 times, then the probability of obtaining at least one head is :
CUET UG 2023 — Mathematics Algebra
If a fair coin is tossed 10 times, then the probability of obtaining at least one head is :
Held on 23 May 2023 · Verified 13 Jul 2026.
Options
- A
10241
- B
102417
- C
10241023
- D
102423
Did you get this right?
Sign in to track your attempts and accuracy.
Your note
Sign in to keep a private note on this question. Nothing you write is ever public.
CUET UG Algebra in other years
More CUET UG Algebra questions
If f(x) = 2x + 3, then f⁻¹(x) is:
If $\vec{a}$ is a unit vector perpendicular to both the vectors $\vec{b} = \hat{j} + \hat{2k}$ and $\vec{c} = \hat{i} + 2\hat{j}$, then $\hat{a}$ is equal to
If the roots of the equation x² - 5x + k = 0 are in the ratio 2:3, then the value of k is:
A function $f: \mathbb{R} \rightarrow \{x \in \mathbb{R}: -1 < x < 1\}$ is defined as $f(x) = \frac{x}{1+|x|}$, then $f$ is:
If A is a square matrix such that $A^2 = A$ and I is the identity matrix of same order as A, then the value of $(A-2I)^2 - (2A + I)^2 + 11A$ is:
Other Mathematics topics for CUET UG
Keep practicing Algebra
Work through every CUET UG Algebra PYQ, year by year.